Circular No. 6072

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET REINMUTH 1 (1994p) Independent recoveries of this comet have been reported by A. Naka-

mura and J. V. Scotti, as follows; the comet is close to the prediction on MPC 20123 (ephemeris on MPC 23873):

1994 UT  α_{2000}  δ_{2000}  m_{1} Observer
Sept.1.68495 2^{h}40^{m}57.^{s}41 +5^{o}15′07′′.0 19.9 Nakamura
1.70087 2 40 57.49 +5 15 02.0  _{”}
1.71834 2 40 57.60 +5 15 01.6  _{”}
3.70955 2 41 11.64 +5 09 55.5 20.1  _{”}
3.76013 2 41 11.89 +5 09 48.0  _{”}
6.41378 2 41 23.38 +5 02 21.9 19.6 Scotti
6.42217 2 41 23.38 +5 02 20.1  _{”}
6.44706 2 41 23.40 +5 02 15.6 20.3  _{”}
6.46881 2 41 23.43 +5 02 11.6 19.8  _{”}

A. Nakamura (Kuma Kogen). 0.60-m f/6.0 Ritchey-Chrétien telescope +

CCD. Comet nearly stellar. Communicated by S. Nakano. J. V. Scotti (Kitt Peak). Spacewatch telescope. 9′′ coma; 0′.30 tail in p.a.

262^{o}; m_{2} = 22.0.

GRO J0422+32 P. Zhao, P. Callanan, M. Garcia, and J. McClintock, Smithsonian As-

trophysical Observatory, write: “The x-ray nova J0422+32 has declined another magnitude since April. Average magnitudes for the period Apr. 12–16 are R = 19.93 ± 0.17 (rms) and V = 20.83 ± 0.13 (rms). CCD photometry obtained with the F. L. Whipple Observatory’s 1.2-m telescope in September gives the following magnitudes: Sept. 5.47 UT, R = 21.06 ± 0.10, V = 22.39 ± 0.27; 7.49, 21.07 ± 0.10, –.”

SUPERNOVA 1994W IN NGC 4041 Visual magnitude estimates (cf. IAUC 6058): Aug. 27.79 UT, 13.7 (L.

Szentaskó, Veresegyház, Hungary); 28.85, 13.6 (G. Cortini and M. Villi, Mount Colombo, Italy); 29.86, 13.4 (W. Hasubick, Buchloe, Germany); 31.79, 14.2 (Szentaskó); Sept. 1.79, 14.3 (Szentaskó); 2.8, 14.3 (Szentaskó).

1994 September 7 Daniel W. E. Green